### v2

@@ Abstract
The ubiquitous noise in quantum system hinders the advancement of quantum information processing and has driven the emergence of different quantum error correction protocols. Among them, quantum error correction codes tailored for noise-biased qubits exhibit comparatively high error thresholds, making them a promising platform to achieve fault-tolerance. Nevertheless, their quantum operations are challenging and the demonstration of their performance beyond the fault-tolerant threshold remains incomplete. Here, we leverage Schrödinger cat states in scalable planar superconducting circuits to thoroughly characterize the high-fidelity single-qubit quantum operations on noise-biased qubits with systematic quantum tomography and benchmarking tools, demonstrating their state-of-the-art performance beyond the fault-tolerant threshold of the surface code. These results thus embody a transformative milestone in the exploration of quantum systems with structured noises. Notably, our framework is extensible to other structured-noise systems, paving the way for systematic characterization and validation of novel quantum platforms with structured noise.
@@
Fault Tolerant, Benchmarking, Noise-biased Qubit, Quantum Error Correction, Cat Qubit
1 Introduction
Rapid advancements in building quantum systems with scalable, well-characterized qubits have opened up the potential to solve problems intractable for classical computers[1, 2, 3]. However, quantum systems remain fragile, subject to various errors such as decoherence[4], dephasing[5, 6], stochastic and coherent errors[7]. As a result, quantum error correction (QEC) is essential for practical and fault-tolerant quantum computing (FTQC)[8, 9]. Most QEC protocols impose stringent performance requirements on the quantum devices due to significant hardware overhead and low error thresholds[9, 10]. The suppression of qubit error by scaling up a surface code reported in the transmon-based device (“Willow”) highlighted the progress in this direction[10]. Alternatively, tailored QEC codes explore the noise structures of quantum systems to achieve higher efficiency and error thresholds, relaxing the requirements for the quantum device performance[11, 12, 13, 14]. QEC codes tailored for noise-biased qubits which are resilient to some of the Pauli noises, such as the repetition code[15] and XZZX surface code[16, 17, 18, 19, 20, 21, 22], have gained extensive attention. Recent work on enhancing the phase-flip lifetime of dissipative cat qubits using repetition codes marks a step toward the extreme limit where qubit noises are so strongly biased that only one error channel requires correction[15]. Nevertheless, universal gate control and quantum nondemolition readout for such qubits remain unclear[23, 24], and additional error channels beyond phase flips are still non-negligible[15]. In contrast, the exploration of Kerr-cat qubits (KCQs) with moderate noise bias has motivated the development of the XZZX surface code, which corrects general stochastic Pauli errors with higher error thresholds. Importantly, KCQs have recently been shown promising through demonstrations of their universal gate control and quantum nondemolition readout[23, 25].
In terms of implementation, the noise bias characteristics can be realized in bosonic qubits by encoding qubits into bosonic modes of quantum oscillators. Representative examples includes the binomial qubits[26], Gottesman-Kitaev-Preskill (GKP) qubits[27, 28] and cat qubits[29, 25, 30, 19, 31, 32, 33, 34, 35, 36, 23, 37]. Among these, cat qubits, achieved by engineering Schrödinger cat states in quantum oscillators, have attracted considerable attention due to their comparatively high noise bias and low hardware complexity. Leveraging strong non-linear interactions in superconducting circuits, many theoretical proposals examined cat qubit implementations[38, 21], bias-preserving gates[39, 40], and applications in XZZX surface code for FTQC[19]. Furthermore, the observation of enhanced bit-flip lifetime of cat qubits encoded in superconducting oscillators has driven experimental advances[30, 25], including the universal gate control of KCQs[23, 41]. However, while the lifetime and gate control of noise-biased qubits have been studied, a comprehensive characterization of their noise structures remains unexplored. The reported gate fidelities are limited to around 80% to 90% due to limited device coherence and state-preparation-and-measurement (SPAM) errors[25, 23], which fall significantly short of the threshold required by many QEC codes[19, 20, 22]. Therefore, thoroughly characterizing the noise structure and obtaining SPAM-free estimations of the gate fidelities are crucial toward effectively applying noise-biased qubits in QEC[42].
In this work, we characterize the noise structure and gate performance of single-qubit operations in planar superconducting circuits for both conventional resonant-KCQ and a novel detuned-KCQ, achieving performance beyond the fault tolerant threshold of the XZZX surface code. Notably, we realize a novel detuned-KCQ with a red-detuned stabilization drive to enhance the bit-flip lifetime. While quantum operations were only reported on conventional resonant-KCQs[43, 44, 45], our results not only showcase high-fidelity universal single-qubit gate control and quantum non-demolition readout for the detuned-KCQ, but also demonstrate that its single-qubit gate performance crosses the fault-tolerance threshold of the XZZX surface code[19]. Remarkably, using gate set tomography (GST), we observe SPAM-free gate fidelities over 92% for the $X(\pi/2)$ gate and 99% for the bias-preserving $Z(\pi/2)$ gate, underscoring state-of-the-art performance. Benchmarking with $\mathbb{D}_{8}$ dihedral group randomized benchmarking (DRB)[46, 42] further reveals the the noise bias exceeding 100 for noise-preserving operations. Importantly, DRB shows that the detuned-KCQ exhibits lower bit-flip errors and similar phase-flip errors compared to the resonant-KCQ, resulting in an even higher noise bias up to 250. With this enhanced noise bias, the single-qubit gate error of the detuned-KCQ goes below the fault-tolerant threshold of the XZZX surface code[19], highlighting the feasibility of QEC with detuned-KCQs. These findings position superconducting Kerr-cat qubits as strong candidates for noise-biased QEC codes, paving the way toward FTQC with noise-biased qubits.
Figure 1: Kerr-cat qubit concepts. a SNAIL nonlinear oscillator circuit with frequency $\omega_{Q}$, nonlinear factors $g_{3},g_{4}$ and Kerr nonlinearity $K$, and its energy levels with wave-functions. b SNAIL oscillator pumped by a two-photon stabilization drive at $\omega_{S}=2\omega_{Q}$, and the engineered pseudo-potential profile with its energy levels and wave-functions. c Pseudo-potential profile and its energy levels engineered by a red-detuned two-photon stabilization drive, where the excited states become degenerate. d Bloch sphere defined by the coherent states and their superpositions, where |±$|\pm Z\rangle=|\pm\alpha\rangle$, |±$|\pm X\rangle=|C_{\alpha}^{\pm}\rangle\propto|\alpha\rangle\pm|-\alpha\rangle$, |±$|\pm Y\rangle=|C_{\alpha}^{\pm i}\rangle\propto|\alpha\rangle\pm i|-\alpha\rangle$. e Pseudo-potential profile under a single-photon drive $\Omega$ at $\omega_{Q}$ for $Z(\theta)$ gates. f Pseudo-potential profile under a blue-detuned two-photon stabilization drive for the $X(\pi/2)$ gate. The detuning was chosen to be $5K$ as an example.
2 Results
2.1 Noise-Biased Cat Qubit Encoding and Control
We implement the KCQ in a superconducting nonlinear oscillator by engineering its effective Hamiltonian under a strong two-photon stabilization drive. The encoding and operations of the KCQ are illustrated in Fig. 1. The nonlinear oscillator consists of two Superconducting Nonlinear Asymmetric Inductive eLements (SNAILs) shunted by a capacitor shown in Fig. 1a. Each SNAIL consists of three large Josephson junctions with Josephson energy $E_{J}$ and a small Josephson junction with Josephson energy $\beta E_{j}$ (Suppl. Info. Section 1). For a fixed external flux $\phi_{\text{ext}}$, the undriven SNAIL nonlinear oscillator offers an anharmonic potential with non-zero third-order $g_{3}$ and fourth-order $g_{4}$ nonlinearities shown in Fig. 1a.
Under a two-photon stabilization drive at around twice the SNAIL oscillator frequency ($\omega_{S}\approx 2\omega_{Q}$), the effective Hamiltonian in the frame rotating at frequency $\omega_{S}/2$ is
$\hat{H}_{\text{KCQ}}/\hbar=\Delta\hat{a}^{\dagger}\hat{a}+\epsilon_{2}\hat{a}^{\dagger 2}+\epsilon_{2}^{*}\hat{a}^{2}-Ka^{\dagger 2}a^{2},$ (1)
where we applied the rotating wave approximation (Suppl. Info. Section 3). The detuning $\Delta=\omega_{Q}-\omega_{S}/2$ is crucial to encode the detuned-KCQ with enhanced performance. The two-photon stabilization drive $\epsilon_{2}$ is engineered with the third-order nonlinearity, which converts a single photon from the drive at frequency $\omega_{S}$ into two photons of the SNAIL oscillator at frequency $\omega_{Q}$, and the Kerr coefficient $K$ is from the fourth-order nonlinearity.
This Hamiltonian forms a pseudo-potential with a double-well structure containing two degenerate ground states at the bottom of the wells, as shown in Fig. 1b and Fig. 1c. These states are separated from the excited states $|\psi_{e}^{\pm}\rangle$ by an energy gap of $4K|\alpha|^{2}$ with $|\alpha|^{2}=(|\epsilon_{2}|+\Delta/2)/K$ characterizing the cat qubit size. These ground states approach the coherent states |±$|\pm\alpha\rangle$ rapidly as $|\epsilon_{2}|/K$ increases (Suppl. Info. Section 3). By encoding these ground states as the $Z$-axis states on the Bloch sphere, their superposition spans the computational space of the KCQ, as shown in Fig. 1d.
As a noise biased qubit, KCQ has higher bit-flip lifetime $T_{z}$ than the phase-flip lifetime $T_{y}$. The coherent states on $Z$ axis, as eigenstates of the bosonic annihilation operator, are immune to photon losses, thus leading to the increased $T_{z}$. However, single-photon loss flips the cat states on the equator of the Bloch sphere with a rate enhanced by $|\alpha|$, leading to the reduced $T_{y}$. It is important to note that, as illustrated in Fig. 1c, the excited states are degenerate when the two-photon stabilization drive is red-detuned by $\Delta=2K$, which leads to an even greater enhancement of the bit-flip lifetime $T_{z}$ (Methods Section 4.1).
Building on our previous work of the single-qubit universal operations on a resonant-KCQ[23], we explored the single-qubit universal quantum operations on a detuned-KCQ through pseudo-potential deformation (Methods Section 4.3). The single-qubit universal gate set generally consists of continuous $Z(\theta)$ gates and an $X(\pi/2)$ gate[47]. On the one hand, by introducing a single photon drive $\Omega$, as shown in Fig. 1e, the energies of two coherent states are shifted, resulting in a continuous $Z(\theta)$ gate. On the other hand, by blue-detuning the stabilization drive, as shown in Fig. 1f, the energy barrier between two coherent states is lowered, allowing them to tunnel between the two wells. Therefore, we can temporarily blue-detune the stabilization drive to turn on such tunneling and achieve a discrete $X(\pi/2)$ gate.
2.2 Qubit Initialization and Lifetime Characterization
The KCQ is initialized to $|+Z\rangle$ state by heralding through cat qubit readout (CQR). CQR is realized by driving the SNAIL oscillator at the frequency difference with the readout resonator to engineer a beam-splitter interaction (Methods Section 4.2)[25, 23]. Under such beam-splitter interaction, the readout resonator is populated to a steady state depending on the average value of the KCQ mode $\langle\hat{a}\rangle$. Therefore, by detecting the readout resonator state, one can determine in which well the KCQ state resides, and collapse it into one of the pseudo-potential wells. This process enables the readout of KCQ along $Z$ axis because the |±$|\pm Z\rangle$ states of KCQ are coherent states localized in one of the potential wells. Owing to the two-photon stabilization drive, such readout scheme is quantum nondemolition as long as the beam splitter interaction strength is much smaller than the two-photon drive strength. We achieved the quantum-nondemolition-ness (QNDness) over 98% which is defined as the probability of measuring the same results in two consecutive readout. Such high QNDness allows the initialization of qubits with such quantum nondemolition readout through heralding.
The initialization sequence is shown in Fig. 2a, along with the corresponding potential profiles and wave-functions in Fig. 2b. By ramping up the two-photon stabilization drive, the potential of the SNAIL oscillator is converted to a double-well pseudo-potential. Notably, due to the detuning $\Delta$ of the two-photon stabilization drive, the vacuum state may not be mapped to the ground state of the pseudo-potential. Therefore, a 400-$\mathcal{\mu}\text{s}$ relaxation time is added to allow the state to decay to the ground states. Finally, the KCQ state is collapsed to the $|+Z\rangle$ states by heralding through CQR (Methods Section 4.3). Fig. 2c shows simulated overlaps of the KCQ state $|\psi(t)\rangle$ with the computational states $|C^{\pm}_{\alpha}\rangle=|\pm X\rangle$ and the vacuum state $|0\rangle$ during initialization with $\Delta=2K$. The leakage out of the computational basis is defined as $p_{\text{leakage}}=1-|\langle\psi(t)|C^{+}_{\alpha}\rangle|^{2}-|\langle\psi(t)|C^{-}_{\alpha}\rangle|^{2}$.
Figure 2: Initialization and lifetimes. a The sequence for KCQ initialization. b The evolution of the potential profiles and wave-functions during the initialization. c The numerical simulation of the KCQ state population transfer during the initialization. d The bit-flip lifetime $T_{z}$ of resonant-KCQ and detuned-KCQ as a function of mean photon numbers. e The phase-flip lifetime $T_{y}$ of resonant-KCQ and detuned-KCQ. The error bars are from fitting uncertainty, and the theory predictions of the lifetimes are calculated by master-equation simulations illustrated in [37] and Methods Section 4.1.
We characterize the bit-flip lifetime $T_{z}$ (phase-flip lifetime $T_{y}$) of the KCQ by initializing it in $Z$-axis ($Y$-axis) and measuring the decay of the Pauli operator expectation values $\langle\hat{\sigma_{z}}\rangle$ ($\langle\hat{\sigma_{y}}\rangle$). As shown in Fig. 2d and Fig. 2e, $T_{z}$ increases quasi-exponentially in a staircase-like manner with the mean photon number $\langle\hat{n}\rangle=|\alpha|^{2}$, while $T_{y}$ decreases according to $T_{y}=T_{1}/(2\langle\hat{n}\rangle)$[23], where $T_{1}\approx 40~{}\text{us}$ is the relaxation time of the SNAIL oscillator. A master equation simulation, accounting for both single- and two-photon losses, photon heating and non-Markovian noise mechanisms, captured the behavior of $T_{z}$, with discrepancies at high $\langle\hat{n}\rangle$ mainly due to increased thermal population (Methods Section 4.1).
Due to the degeneracy of higher excited states, we observed increased $T_{z}$ with detuning $\Delta=2K,4K...$ (Methods Section 4.1). Specifically, $T_{z}$ is higher for $\Delta=2K$ compared to $\Delta=0$ with the same mean photon number, and more importantly, the same $T_{y}$, as shown in Fig. 2d and 2e. Therefore, a detuned-KCQ with $\Delta=2K$ is expected to exhibit greater noise bias without additional phase-flip errors, which motivates the exploration of detuned-KCQ for enhanced performance[43, 44].
@@ 2.2 Qubit Initialization and Lifetime Characterization
2.3 Gate Fidelity Characterization
@@
To estimate the performance of gate operations on resonant-KCQ and detuned-KCQ, we implement GST to extract the error rates, gate fidelities, and SPAM-free Pauli transfer matrices (PTMs) (Suppl. Info. Section 5). The error channels are considered as a combination of coherent and stochastic Pauli errors, described by the PTM $\mathcal{E}=e^{\mathbf{L}}$. The error generator $\mathbf{L}$ is defined as
@@
where $\mathbf{H}_{x}(\mathbf{P}_{x}),\mathbf{H}_{y}(\mathbf{P}_{y}),\mathbf{H}_{z}(\mathbf{P}_{z})$ are the coherent (stochastic) error generators with respect to $X$, $Y$ and $Z$ axis and $h_{x}~{}(p_{x}),h_{y}~{}(p_{y}),h_{z}~{}(p_{z})$ are the corresponding errors (Suppl. Info. Section 7).
The model violation of GST under completely positive, trace preserving (CPTP) constraints is as small as $N_{\sigma}=14.51$, indicating the validity of the model. The SPAM-free estimation of the PTMs, errors and fidelities for $X(\pi/2)$ and $Z(\pi/2)$ gates are shown in Fig. 3. For $Z(\pi/2)$ gates, the stochastic Pauli $X$ and $Y$ errors are significantly smaller than the stochastic Pauli $Z$ errors, confirming the bias preservation. Our KCQs have low coherent errors, as indicated by the similar diamond distances and process infidelity. In fact, the effects of coherent errors are negligible after Pauli twirling and have no statistically observable effects on the XZZX surface code error threshold (Suppl. Info. Section 7). According to the extracted process infidelities of the quantum gates, the process fidelities of $Z(\pi/2)$ and $X(\pi/2)$ gates for resonant-KCQ are 99.3% and 94.7%, and the detuned-KCQ has similar performance with the fidelities of 99.2% and 92.5%, respectively, signifying state-of-the-art performance[25, 23].

### v3

@@ Abstract
Ubiquitous noises in quantum systems remain a key obstacle to building quantum computers, necessitating the use of quantum error correction codes. Recently, error-correcting codes tailored for noise-biased systems have been shown to offer high fault-tolerance thresholds and reduced hardware overhead, positioning noise-biased qubits as promising candidates for building universal quantum computers. However, quantum operations on these platforms remain challenging, and their noise structures have not yet been rigorously benchmarked to the same extent as those of conventional quantum hardware. In this work, we develop a comprehensive quantum control toolbox for a scalable noise-biased qubit, detuned Kerr-cat qubit, including initialization, universal single-qubit gates and quantum non-demolition readout. We systematically characterize the noise structure of these operations using gate set tomography and dihedral randomized benchmarking, achieving high local gate fidelities, with $\mathcal{F}[Z(\pi/2)]=99.2\%$ and $\mathcal{F}[X(\pi/2)]=92.5\%$. Notably, the noise bias of the detuned Kerr-cat qubit approaches 250, which outperforms its resonant-Kerr-cat qubit counterparts as reported previously, representing a new state-of-the-art performance benchmark for noise-biased qubits. Moreover, our results reveal a critical overestimation of operational noise bias inferred from bit-flip and phase-flip times alone, highlighting the necessity of a precise and direct benchmarking for noise-biased qubit operations. Our work thus establishes a framework for systematically characterizing and validating the performance of quantum operations in structured-noise architectures, which lays the groundwork for implementing efficient quantum error correction in next-generation architectures.
@@
noise-biased qubit $|$ quantum benchmarking $|$ quantum error correction
@@
This manuscript was compiled on August 8, 2025 www.pnas.org/cgi/doi/10.1073/pnas.XXXXXXXXXX
4]2
[…]
Rapid advancements in building quantum systems with scalable, well-characterized qubits have opened up the potential to solve problems intractable for classical computers (1, 2, 3). However, quantum systems remain fragile, subject to various errors such as decoherence (4), dephasing (5, 6), stochastic and coherent errors (7). As a result, quantum error correction (QEC) is essential for practical and fault-tolerant quantum computing (FTQC) (8, 9). Most QEC protocols impose stringent performance requirements on the physical device level due to significant hardware overhead and low error thresholds (9, 10). The suppression of qubit error via surface code highlighted the progress in this direction (10). Alternatively, tailored QEC codes explore the noise structures of quantum systems to achieve higher efficiency and error thresholds, relaxing the requirements for the quantum device performance (11, 12, 13, 14). QEC codes tailored for noise-biased qubits (15, 16, 17), which are resilient to some of the Pauli noises, such as the XZZX codes (18, 19, 20), have gained extensive attention. Recent work on enhancing the phase-flip time of dissipative cat qubits using repetition codes marks an important step toward the extreme limit where physical noises are so strongly biased that only one error channel requires correction (21). Nevertheless, the additional error channels beyond phase flips are still non-negligible (21). On the other hand, the exploration of Kerr-cat qubits (KCQs) aims to address stochastic Pauli errors based on the XZZX surface code and relies on noise bias to achieve high error thresholds (22). Importantly, resonant-KCQs have recently been shown to be promising through demonstrations of their universal gate control and quantum nondemolition readout (23, 24).
In terms of implementation, noise bias can be achieved by encoding qubits into bosonic modes of quantum oscillators. Specifically, cat qubits (25, 24, 26, 22, 27, 28, 29, 30, 31, 32, 23, 33), realized by engineering Schrödinger cat states in quantum oscillators, have attracted considerable attention due to their comparatively high noise bias and low hardware complexity. Leveraging strong non-linear interactions in superconducting circuits, many theoretical proposals examined cat qubit implementations (34, 19), bias-preserving gates (35, 36), and applications in XZZX surface code for FTQC (22). Furthermore, the successful implementation of cat qubits in superconducting circuits has driven experimental advances  (26, 24). The universal quantum operations of a KCQ under a resonant two-photon stabilization drive, or resonant-KCQ, have been demonstrated recently (23, 37, 38). In addition, a significant bit-flip time improvement has been observed on detuned-KCQs through a red-detuned two-photon stabilization drive (39). However, noise structure characterization on KCQs has so far been limited to bit-flip and phase-flip time measurements when the qubits are idle, leading to overestimated noise bias in actual gate operations. Furthermore, while detuned-KCQs have improved bit-flip times (39), their universal quantum operations have not yet been demonstrated. The absence of direct noise characterization for KCQ operations and universal gate implementations for detuned-KCQs has limited their further progress. Therefore, the development of a comprehensive quantum control toolbox for detuned-KCQs and a direct, accurate, and thorough characterization of the quantum gate noise structures are crucial for advancing KCQs towards further applications in QEC.
Figure 1: Kerr-cat qubit concepts. a, SNAIL nonlinear oscillator circuit with frequency $\omega_{Q}$, nonlinear factors $g_{3},g_{4}$ and Kerr nonlinearity $K$, and its energy levels with wave-functions. b, SNAIL oscillator pumped by a two-photon stabilization drive at $\omega_{S}=2\omega_{Q}$, and the engineered pseudo-potential profile with its energy levels and wave-functions. c, Pseudo-potential profile and its energy levels engineered by a red-detuned two-photon stabilization drive, where the excited states become degenerate. d, Bloch sphere defined by the coherent states and their superpositions, where $|\pm Z\rangle=|\pm\alpha\rangle$, $|\pm X\rangle=|C_{\alpha}^{\pm}\rangle\propto|\alpha\rangle\pm|-\alpha\rangle$, $|\pm Y\rangle=|C_{\alpha}^{\pm i}\rangle\propto|\alpha\rangle\pm i|-\alpha\rangle$. e, Pseudo-potential profile under a single-photon drive $\Omega$ at $\omega_{S}/2$ for $Z(\theta)$ gates. f, Pseudo-potential profile under a blue-detuned two-photon stabilization drive for the $X(\pi/2)$ gate. The detuning was chosen to be $5K$ as an example.
In this work, we develop the first comprehensive quantum control toolbox for the detuned-KCQ in planar superconducting circuits, including the high-fidelity initialization, universal single-qubit gates and quantum nondemolition readout, with a comparison to its resonant counterpart realized in our previous work (23). Moreover, we directly characterize the noise structures of single-qubit operations for both the resonant- and detuned-KCQ, providing a precise and detailed description of the various noises affecting these operations. Using gate set tomography (GST), we achieved SPAM-free gate fidelities over 92.5% for the $X(\pi/2)$ gate and 99.2% for the bias-preserving $Z(\pi/2)$ gate, underscoring the highest performance in a noise-biased platform. Benchmarking with $\mathbb{D}_{8}$ dihedral group randomized benchmarking (DRB) (40, 41) further reveals the noise bias exceeding 100 for their noise-preserving operations. Notably, the detuned-KCQ exhibits lower bit-flip errors and similar phase-flip errors compared to the resonant-KCQ, resulting in a noise bias of 250.
Interestingly, the noise bias of short gates extracted using DRB is much lower than that estimated from the bit-flip and phase-flip time measurements, signifying the inaccuracy of noise bias characterization based solely on idling lifetimes. As the first thorough and direct exploration of the noise structures of quantum operations on noise-biased qubits, we provide a framework for benchmarking the performance of quantum systems with structured noises. Moreover, thanks to the development of high-fidelity universal single-qubit operations on the detuned-KCQ, our work also positions detuned-KCQ as a strong candidate for noise-biased QEC.
Noise-Biased Cat Qubit Encoding and Control
We extend the investigation of the resonant-KCQ introduced in our previous work (23) to explore the detuned-KCQ by engineering the effective Hamiltonian of a superconducting nonlinear oscillator under a strong two-photon stabilization drive with various frequencies. The encoding and operations of the detuned-KCQ are illustrated in Fig. 1. The nonlinear oscillator consists of two Superconducting Nonlinear Asymmetric Inductive eLements (SNAILs) shunted by a capacitor shown in Fig. 1a. Each SNAIL consists of three large Josephson junctions with Josephson energy $E_{J}$ and a small Josephson junction with Josephson energy $\beta E_{j}$ (Suppl. Info. Section 1). For a fixed external flux $\phi_{\text{ext}}$, the undriven SNAIL nonlinear oscillator offers an anharmonic potential with non-zero third-order $g_{3}$ and fourth-order $g_{4}$ nonlinearities shown in Fig. 1a.
Under a two-photon stabilization drive at frequency $\omega_{S}$, the effective Hamiltonian in the frame rotating at frequency $\omega_{S}/2$ is
$\hat{H}_{\text{detuned-KCQ}}/\hbar=\Delta\hat{a}^{\dagger}\hat{a}+\epsilon_{2}\hat{a}^{\dagger 2}+\epsilon_{2}^{*}\hat{a}^{2}-K\hat{a}^{\dagger 2}\hat{a}^{2},$ (1)
where we applied the rotating wave approximation (Materials and Methods). The non-zero detuning $\Delta=\omega_{Q}-\omega_{S}/2$ is crucial to encode the detuned-KCQ with enhanced performance. The two-photon stabilization drive $\epsilon_{2}$ is engineered with the third-order nonlinearity, which converts a single photon from the drive at frequency $\omega_{S}$ into two photons in the SNAIL oscillator at frequency $\omega_{Q}$, and the Kerr coefficient $K$ originates from the fourth-order nonlinearity.
This Hamiltonian forms a pseudo-potential with a double-well structure containing two degenerate ground states at the bottom of the wells, as shown in Fig. 1b for resonant-KCQ with $\Delta=0$ and Fig. 1c for detuned-KCQ with $\Delta=2K$. These states are separated from the excited states $|\psi_{e}^{\pm}\rangle$ by an energy gap of $4K|\alpha|^{2}$ with $|\alpha|^{2}=(|\epsilon_{2}|+\Delta/2)/K$, where $|\alpha|^{2}$ is often used to indicate the size of the cat in phase space. These ground states approach the coherent states $|\pm\alpha\rangle$ rapidly as $|\epsilon_{2}|/K$ increases (Materials and Methods). By encoding these ground states as the $Z$-axis states on the Bloch sphere, their superposition spans the computational space of KCQs, as shown in Fig. 1d.
As a noise-biased qubit, KCQ has a higher bit-flip time $T_{z}$ than the phase-flip time $T_{y}$. The coherent states on the $Z$ axis, as eigenstates of the bosonic annihilation operator, are immune to photon losses, thus leading to the increased $T_{z}$. However, single-photon loss flips the cat states on the equator of the Bloch sphere with a rate enhanced by $|\alpha|$, leading to the reduced $T_{y}$. It is important to note that, as illustrated in Fig. 1c, the excited states of the detuned-KCQ are degenerate when the two-photon stabilization drive is red-detuned by $\Delta=2K$. Such degeneracy can lead to an even higher bit-flip time $T_{z}$ without further degradation of the phase-flip time $T_{y}$ (39), providing advantages in quantum information processing (Materials and Methods).
Figure 2: Initialization, bit-flip times and phase-flip times. a, The sequence for detuned-KCQ initialization. b, The evolution of the potential profiles and wave-functions during the initialization. c, The numerical simulation of the detuned-KCQ state population transfer during the initialization. d, The bit-flip time $T_{z}$ of resonant-KCQ and detuned-KCQ as a function of mean photon numbers. e, The phase-flip time $T_{y}$ of resonant-KCQ and detuned-KCQ. The error bars are from fitting uncertainty, and the theory predictions of the bit-flip times are calculated by master-equation simulations illustrated in  (33, 39) and Materials and Methods.
While the quantum operations of a resonant-KCQ have been extensively studied in our previous work and related literature (23, 24, 38), only bit-flip time measurement has been reported for detuned-KCQs with a small two-photon stabilization drive (39), hindering its further development. Here, we demonstrate a complete quantum control toolbox for detuned-KCQ, including the initialization, universal single-qubit gates, and quantum nondemolition readout. The universal single-qubit gates consist of the continuous $Z(\theta)$ gates and a discrete $X(\pi/2)$ gate through pseudo-potential deformation. The continuous $Z(\theta)$ gates are realized by a single-photon drive $\Omega$ to lift the degeneracy between the two coherent states, allowing them to accumulate different dynamical phases, as shown in Fig. 1e. The discrete $X(\pi/2)$ gate is engineered by blue-detuning the stabilization drive to lower the energy barrier between the two coherent states, allowing them to tunnel between the two wells, as shown in Fig. 1f (Materials and Methods). In contrast to the previously reported method for resonant-KCQs based on the interruption of two-photon stabilization drive (38), our method corresponds to a fidelity over 92%, even for large cat sizes, by keeping the stabilization drive on to mitigate the effects of decays of Fock states.
Qubit Initialization, readout and Lifetime Characterization
The KCQ is initialized to the $|+Z\rangle$ state by heralding through cat quadrature readout (CQR). CQR is realized by a beam-splitter interaction engineered through the three-wave mixing process (Materials and Methods). Under such a beam-splitter interaction, the readout resonator is populated to a steady state depending on the average value of the KCQ mode $\langle\hat{a}\rangle$. Therefore, by detecting the readout resonator state, one can determine in which well the KCQ state resides and collapse it into one of the pseudo-potential wells. This process enables the readout of KCQ along the $Z$ axis because the $|\pm Z\rangle$ states of KCQ are coherent states localized in one of the potential wells. This readout method is quantum nondemolition when the mean photon numbers in KCQs are conserved due to the much stronger two-photon stabilization drives than the beam splitter interactions. The quantum-nondemolition-ness (QNDness), defined as the probability of measuring the same results in two consecutive readouts, is observed to be over 98%. Such high QNDness allows the initialization of qubits through heralding.
While the resonant-KCQ can be initialized from a vacuum state by applying a CQR pulse followed immediately after ramping up the stabilization drive, a sufficiently long relaxation time before the CQR pulse is necessary to initialize the detuned-KCQ. This is because the vacuum state can be mapped to the excited states with a detuned stabilization drive. (Materials and Methods). In the experiments, the initialization sequence is shown in Fig. 2a, along with the corresponding potential profiles and wave-functions in Fig. 2b. By ramping up the two-photon stabilization drive, the potential of the SNAIL oscillator is converted to a double-well pseudo-potential. Notably, due to the detuning $\Delta$ of the two-photon stabilization drive, the vacuum state is mapped to the even-parity excited state of the pseudo-potential, which is out of the computational manifold. Therefore, a 400-$\mathcal{\mu}\text{s}$ relaxation time is added to allow this excited state to decay to the ground states in the computational manifold. Finally, the detuned-KCQ state is collapsed onto the $|+Z\rangle$ states by a CQR pulse (Materials and Methods).
Fig. 2c shows the simulated overlaps of the detuned-KCQ state $|\psi(t)\rangle$ with the computational states $|C^{\pm}_{\alpha}\rangle=|\pm X\rangle$ and the vacuum state $|0\rangle$ during initialization with $\Delta=2K$. The leakage out of the computational basis is defined as $p_{\text{leakage}}=1-|\langle\psi(t)|C^{+}_{\alpha}\rangle|^{2}-|\langle\psi(t)|C^{-}_{\alpha}\rangle|^{2}$. Another fast initialization method based on a chirp pulse has been reported recently (37, 42), but our method achieves higher initialization fidelity despite the long relaxation times, as indicated by the high QNDness over 98% (Materials and Methods, Fig. 7).
[…]
Due to the degeneracy of higher excited states, we observed increased $T_{z}$ with detuning $\Delta=2K,4K$... (Materials and Methods). Specifically, $T_{z}$ is higher for $\Delta=2K$ compared to $\Delta=0$ with the same mean photon number, and more importantly, the same $T_{y}$, as shown in Fig. 2d and 2e. Therefore, a detuned-KCQ with $\Delta=2K$ is expected to exhibit greater noise bias without additional phase-flip errors, which motivates the exploration of detuned-KCQ for enhanced performance (39, 45).
[…]
Gate Fidelity Characterization
@@
To estimate the performance of gate operations and, more importantly, the noise structures on resonant- and detuned-KCQ, we implement GST to extract the error rates, gate fidelities, and Pauli transfer matrices (PTMs) without SPAM errors (Suppl. Info. Section 5). The error channels are considered a combination of coherent and stochastic Pauli errors, described by a PTM $\mathcal{E}=e^{\mathbf{L}}$. The error generator $\mathbf{L}$ is defined as
@@
where $\mathbf{H}_{x}(\mathbf{P}_{x}),\mathbf{H}_{y}(\mathbf{P}_{y}),\mathbf{H}_{z}(\mathbf{P}_{z})$ are the coherent (stochastic) error generators with respect to $X$, $Y$ and $Z$ axis and $h_{x}\leavevmode\nobreak\ (p_{x}),h_{y}\leavevmode\nobreak\ (p_{y}),h_{z}\leavevmode\nobreak\ (p_{z})$ are the corresponding errors (Suppl. Info. Section 7).